Circular No. 8192

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 QX_{29} (NEAT) An apparently asteroidal object reported by NEAT (Palomar discovery

observation below; originally posted on the NEO Confirmation Page, then assigned the designation 2003 QX_{29} on MPEC 2003-Q33; observations on MPS 93475–93476) has been found to have cometary appearance on CCD images taken by I. Griffin and S. G. Huerta (Cerro Tololo 0.9-m reflector, Aug. 31.1 UT; visible coma of red mag 18.0–19.4 with FHWM = 2′′.3–2′′.6 in raw 300-s images, while stacked 10-exposure image shows a fan-shaped tail at least 17′′ long in p.a. 58^{o}) and by J. Young (Table Mountain 0.6-m reflector, Sept. 1.2; 3′′ coma, slightly elongated in p.a. 260^{o}, with a 16′′ curved tail starting in p.a. 243^{o}; possible slight brightening in the tail at a point 4′′-5′′ from the coma edge). J. Tichá subsequently reports that Kleť images from Aug. 23.9 show the object to be slightly diffuse, while on Aug. 24.9 it exhibited a 8′′ coma. Astrometry, orbital elements (T = 2002 Oct. 17.3 TT, ω = 37^{o}.1, Ω = 264^{o}.9, i = 11^{o}.4, equinox 2000.0, e = 0.445, q = 4.311 AU, P = 21.6 yr), and an ephemeris appear on MPEC 2003-R14.

2003 UT  α_{2000}  δ_{2000}  Mag.
Aug. 23.28243  22^{h}39^{m}04.^{s}71  +6^{o}41′07′′.7  19.8

SUPERNOVA 2003ht IN UGC 2457 Further to IAUC 8191, M. Papenkova and W. Li report the LOTOSS

discovery of an apparent supernova (mag ∼ 18.5) on unfiltered KAIT images on Sept. 1.5 and 2.5 UT. SN 2003ht is located at α = 2^{h}59^{m}55^{s}.17, δ = +24^{o}13′36′′.8 (equinox 2000.0), which is 4′′.3 east and 5′′.0 north of the nucleus of UGC 2457. A KAIT image taken on Jan. 18.2 (limiting mag ∼ 19.5) and 29.2 (limiting mag ∼ 19.0) showed nothing at this position.

SUPERNOVA 2003hq IN MCG +08-30-27 R. Chornock, S. Jha, and A. V. Filippenko, University of California

at Berkeley, report that inspection of CCD spectra (range 320–1000 nm), obtained at the Shane 3-m telescope at Lick Observatory on Sept. 2.2 UT, reveals that SN 2003hq (IAUC 8191) is a type-Ia supernova with a spectralfeature age of 4 (± 2) days before maximum light. The minimum of the strong Si II 635.5-nm absorption feature is observed at 640 nm, implying a host recession velocity of ≈ 12000 km/s (assuming a normal supernova expansion velocity of 10000 km/s).

2003 September 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green